OpenStudy (anonymous):

Evaluate 2x2 − 5y when x = -3 and y = 4. A. -38 B. -32 C. -8 D. -2

OpenStudy (anonymous):

Lets See Whos Smart !

OpenStudy (anonymous):

there's nothing to get confused of... just substitute: \[\LARGE 2\cdot (-3)^2-5\cdot 4\] go on...

OpenStudy (anonymous):

2*(-3)^2 - 5 * 4 =18 - 20 =-2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

why is it all the answers to all your questions r D lol